# Distribution Transformers: Fast Approximate Bayesian Inference With On-The-Fly Prior Adaptation

### v1

@@ 4.1 Analytical Verification Study
In specific cases, where the adopted prior is of a conjugate family to the likelihood, the posterior is tractable. We first verify that our approach does indeed perform approximate inference, for the case of an inverse-gamma prior and normal-variance likelihood. We choose a meta-prior consisting of independent inverse-gamma distributions over the rate and shape parameters, and test our approach on both narrow and wide parameter settings.
@@
Table 1: Results for Experiment 4.1 (best), tested on and timed over 1000 sampled unseen problems. Expected KL-Divergences with the true posterior (along with 95% confidence intervals) are given for the narrow (above) and wide (below) meta-priors. DT-$k$ refers to a $k$-component Distribution Transformer. Note first that both variants of our proposal achieve better posterior KL-Divergences than an equivalent PFN and SVI for both meta-prior settings, while performing inference orders of magnitude faster than SVI. This difference is particularly apparent for the wide meta-prior, where the PFN fails to fit the posterior entirely.
@@
0.0004$\pm$0.0000
0.0003$\pm$0.0000
0.0156
Figure 3 demonstrates that for both wide and narrow meta-priors, the Distribution Transformer does indeed learn to perform Bayesian inference and provides excellent approximations to the posterior, even under change of prior. Figure 1 shows the GMM approximations provided by the DT for this study, demonstrating that even with no additional model parameters a high-quality mapping to a GMM approximation of the prior is achieved. It is clear that while PFNs perform as well as the Riemann distribution allows for the fixed-prior case, as expected in the case of variable priors they fail to fit the posterior at all. This is further demonstrated in Table 1, which shows that not only do we achieve speeds close to that of PFNs, which are slightly faster due to the lighter-weight architecture, but even significantly outperform the much slower SVI in terms of posterior KL-Divergence. These performance gains can be attributed to the expressive GMM adopted by our method, and the efficient transformer-based architecture. An interesting observation, is the extremely high uncertainty in the estimate for the PFN posterior KL-Divergence, marked *. This can be attributed to a failing of the Riemann distribution in this setting—the half-Gaussian tail adopted by the Riemann distribution has variance fitted to the (marginal) prior, meaning for certain observations (or priors), the true posterior has significant probability mass in the right tail which the Riemann distribution cannot express, leading to a skewed distribution for the expected KL-Divergence with a misleading 95%-confidence interval.
@@ 4.2.1 Gaussian Process Joint Predictive Posterior and Hyperposterior
In Table 2 we show that we outperform existing methods in terms of NLL for both PPD and hyperposterior, as expected, while also being the fastest method in terms of runtime. In Figure 4, we show an example PPD for for one set of datapoints and are clearly closer to the oracle PPD that has the knowledge of the true lengthscale.
Figure 4: Example plot for the GP predictive experiment with hyperpriors, using 10 GMM components. Here we put an InverseGamma(1,2) prior on the lengthscale. We plot our model’s predictive posterior in blue and PFNs in green. In orange, we show the oracle that is the exact GP fit with the true lengthscale value (which is unobserved for the other methods). We see that PFNs overestimate the confidence intervals due to the fact that they do not take the prior, particularly that of the lengthscale, into account.The Riemann distribution of the PFN uses 30 buckets.
Table 2: Expected NLL for the marginal PPD and the marginal hyperposterior (denoted Hyper), and inference time per 1000 problems (denoted Runtime), for Experiment 4.2.1. VI is not evaluated for the PPD, as is convention. As expected, we outperform in both categories. Note that PFNs suffer the same issue with the hyperposterior NLL confidence here as the one in Experiment 4.1.
Method Expected NLL
PPD Hyper Runtime (s)
VI 1.97 $\pm$ 0.53 131
PFN 0.00 $\pm$ 0.10 3.21 $\pm$ 4.40 1.03
DT -0.32 $\pm$ 0.12 0.81 $\pm$ 0.08 0.96
@@ 4.2.2 Quantum System Parameter Inference
Figure 5: Expected negative log-likelihood against batch inference time for SVI, a PFN and a 5-component DT. Note that even given orders of magnitude more computation time, SVI cannot match the performance of our method, again demonstrating the power of our GMM approximation. Note that this problem is particularly challenging for VI, as the likelihood must be marginalized with respect to the uncertainty in the initial state and measurement time, which is not tractable and must be estimated stochastically, increasing the time per iteration.
Figure 5 illustrates the power of our approach, achieving better log-likelihood performance (and therefore a smaller KL-Divergence with the true posterior, indicating a better fit), than both competitor methods. This is a problem setting where variable priors are important, as they can be subjective in this setting, and prior sensitivity analysis can be important. To enable such an analysis, fast inference is also important. This illustrates the practical advantage of our approach over PFNs and VI well, as we explicitly target these problems.
@@ 4.3 Sequential Inference Study
The industry standard for real-time sequential inference (or Bayesian filtering), and therefore sensor fusion, is the Extended Kalman Filter (EKF). The EKF linearizes system dynamics and observation models, and approximates all sources of uncertainty as Gaussian. However, these assumptions are rarely reflected in reality. We tackle a sensor fusion problem consisting of 2-dimensional linear dynamics, modelled as a 4-dimensional state space (displacement and velocity), with indirect measurements of displacement provided by two independent, non-linear, non-Gaussian, sensors. Our approach is likely to be successful here, as prior time spent training is almost irrelevant, and fast inference with variable priors is the priority. We also validate against a particle filter, which handles these non-ideal conditions well at the expense of computation time.
@@
Table 3: Expected NLL and iteration time (prediction, update and distribution generation) for EKF, PF, and DT. Note that our method vastly outperforms the EKF in terms of NLL with a small cost in iteration time, while almost matches the close to ground-truth PF’s NLL while achieving close to 50$\times$ speedup.
@@ 5 Conclusion
In summary, Distribution Transformers (DTs) introduce a powerful framework for approximate Bayesian inference by combining universally-approximating Gaussian Mixture Models with Transformer architectures. Our empirical results demonstrate that DTs not only achieve superior inference accuracy and speed compared to existing methods, but also enable dynamic prior updates without retraining—a unique capability in the field. Furthermore, DTs show near-exact inference performance on challenging tasks including GP hyperposterior estimation, quantum parameter inference, and real-world sequential inference, while maintaining the computational efficiency needed for practical applications.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix B General Architecture Details
A standard transformer decoder setup is used, always consisting of 6 transformer decoder units, typically with a latent space size of 64, an MLP hidden layer size of 2048, and 8 attention heads.
@@ Appendix C General Training Details
Table 4: Training hyperparameters and statistics for all experiments. An epoch always consists of 100 batches. Hyperparameters were obtained by limited manual tuning, adjusted until training was stable in all cases. For SVI, the total samples reported account for batching, so the number of samples used by each SVI problem will be a factor of the number of test problems lesser. Parameters are listed vertically, corresponding to DTs (top), PFNs (middle), and SVI (bottom). For Experiment 4.1, only DT-5 is reported but training hyperparameters are identical for DT-2. For experiment 4.3 only DT is reported. For experiment 4.2.1, two PFNs are trained (one for each dimension, as PFNs do not trivially support multivariate distributions), so two figures are quoted when necessary (PPD + hyperposterior). Model sizes are not provided for VI as this is negligible.
@@
For the meta prior in this experiment, we assign a narrow, uniform prior over the PPD’s prior such that it is marginally everywhere a standard normal. We assign a uniform prior on $[1,2]$ to both parameters of the inverse-gamma prior over lengthscale.
@@
The sampling procedure for this experiment, using 5 observations, was as follows:
@@
    Sample $x\sim\text{Uniform}(0,10)$
@@
    Sample $X\sim\text{Uniform}(0,10)^{5}$
@@
For this experiment, we adopt a beta prior, with an $\text{InverseGamma}(4,6)$ meta-prior over each parameter.

### v2

@@ 4.1 Analytical Verification Study
In specific cases, where the adopted prior is of a conjugate family to the likelihood, the posterior is tractable. We first verify that our approach indeed performs approximate inference, for the case of an inverse-gamma prior and normal-variance likelihood. We choose a meta-prior consisting of independent inverse-gamma distributions over the rate and shape parameters, and test our approach on both narrow and wide settings.
@@
Table 1: Results for Experiment 4.1 (best), tested on and timed over 1000 sampled unseen problems. Expected KL-Divergences with the true posterior (along with 95% confidence intervals) are given for the narrow (above) and wide (below) meta-priors. DT-$k$ refers to a $k$-component Distribution Transformer, and PFN-$n$ refers to a PFN equipped with an $n$-bucket Riemann distribution. Note first that both variants of our proposal achieve better posterior KL-Divergences than an equivalent PFN and SVI for both meta-prior settings, while performing inference orders of magnitude faster than SVI. This difference is particularly apparent for the wide meta-prior, where the PFN fails to fit the posterior entirely.
@@
[…]
0.016
Figure 3 demonstrates that for both wide and narrow meta-priors, the Distribution Transformer does indeed learn to perform Bayesian inference and provides excellent approximations to the posterior, even under change of prior. Figure 1 shows the GMM approximations provided by the DT for this study, demonstrating that even with no additional model parameters a high-quality mapping to a GMM approximation of the prior is achieved. It is clear that while PFNs perform as well as the Riemann distribution allows for the narrow meta-prior case, as expected in the wide case they fail to fit the posterior at all. This is further demonstrated in Table 1, which shows that not only do we achieve speeds close to that of PFNs, which are slightly faster due to the lighter-weight architecture, but even significantly outperform the much slower SVI in terms of posterior KL-Divergence. These performance gains can be attributed to the expressive GMM adopted by our method, and the efficient transformer-based architecture. TabPFNv2, while being much bigger of a model than any other baseline, is also unable to improve much on top of standard PFNs. ACE achieves second best performance, losing only to DTs. It thus appears that all methods utilising GMMs surpass those based on Riemannian histogram. However, DTs can provide greater improvement, presumably because self-attention is applied to mixture components (as opposed to data in ACE) allowing the model to spend more computations on shaping the correct distribution shape.
An interesting observation, is the extremely high uncertainty in the estimate for the PFN posterior KL-Divergence, marked *. This can be attributed to a failing of the Riemann distribution in this setting—the half-Gaussian tail adopted by the Riemann distribution has variance fitted to the (marginal) prior, meaning for certain observations (or priors), the true posterior has significant probability mass in the right tail which the Riemann distribution cannot express, leading to a skewed distribution for the expected KL-Divergence with a misleading 95%-confidence interval.
@@ 4.2.1 Gaussian Process Joint Predictive Posterior and Hyperposterior
In Table 2 we show that we outperform existing methods on a more challenging 5-dimensional input problem in terms of NLL for both PPD and hyperposterior, as expected, while also being the fastest method in terms of runtime. In Figure 4, we show an example PPD for for one set of datapoints and are clearly closer to the oracle PPD knowing the true lengthscale.
Figure 4: Example plot for the 1-dimensional input GP predictive experiment with hyperpriors, using 10 GMM components. Here we put an InverseGamma(1,2) prior on the lengthscale. We plot our model’s predictive posterior in blue and PFNs in green. In orange, we show the oracle that is the exact GP fit with the true lengthscale value (which is unobserved for the other methods). We see that PFNs overestimate the confidence intervals due to the fact that they do not take the prior, particularly that of the lengthscale, into account.The Riemann distribution of the PFN uses 30 buckets, matching the number of outputs of the DT.
Table 2: Expected NLL for the marginal PPD and the marginal hyperposterior (denoted Hyper), and inference time per 1000 problems (denoted Runtime), for Experiment 4.2.1. VI is not evaluated for the PPD, as is convention. As expected, we outperform in both categories. Note that PFNs suffer the same issue with the hyperposterior NLL confidence here as in Experiment 4.1. The usually-accurate MCMC also underperforms here, as certain meta-prior settings give rise to instability and non-convergence, and thus inaccuracies.
Method Expected NLL Inference
PPD Hyper Time (s)
VI 0.39 $\pm$ 0.05 123
MCMC 2.27 $\pm$ 0.04 27
PFN 0.90 $\pm$ 0.02 1.53 $\pm$ 0.01 9.2
tabPFNv2 1.08 $\pm$ 0.02 0.37 $\pm$ 0.02 28.5
ACE 1.08 $\pm$ 0.02 0.35 $\pm$ 0.02 11.7
DT 0.81 $\pm$ 0.02 0.31 $\pm$ 0.02 9.5
@@ 4.2.2 Quantum System Parameter Inference
Figure 5: Expected negative log-likelihood against batch inference time. Note that even given orders of magnitude more computation time, SVI cannot match the performance of our method, again demonstrating the power of our GMM approximation. Note that this problem is particularly challenging for VI, as the likelihood must be marginalized with respect to the uncertainty in the initial state and measurement time, which is not tractable and must be estimated stochastically, increasing the time per iteration.
Figure 5 illustrates the power of our approach, achieving better log-likelihood performance (and therefore a smaller KL-Divergence with the true posterior, indicating a better fit), than PFNs and SVI and matching performance of TabPFNv2 and ACE, while being faster. This is a problem setting where variable priors are important, as the prior sensitivity analysis is typically necessary. To enable such an analysis, fast inference is crucial. This further illustrates the practical advantage of our approach, as we explicitly target these problems.
@@ 4.3 Sequential Inference Study
The industry standard for real-time sequential inference (or Bayesian filtering), and therefore sensor fusion, is the Extended Kalman Filter (EKF). The EKF linearizes system dynamics and observation models, and approximates all sources of uncertainty as Gaussian. However, these assumptions are rarely reflected in reality. We tackle a sensor fusion problem consisting of 2-dimensional linear dynamics, modelled as a 4-dimensional state space, with indirect measurements of displacement provided by two independent, non-linear, non-Gaussian, sensors. Our approach is well-suited to this problem setting, as time spent training is almost irrelevant, and fast, conjugate inference with variable priors is the priority. We also validate against a particle filter, which handles these non-ideal conditions well at the expense of computation time.
@@
[…]
@@ 5 Conclusion
In summary, Distribution Transformers (DTs) introduce a powerful framework for approximate Bayesian inference by combining universally-approximating GMMs with Transformer architectures. Our empirical results demonstrate that DTs not only achieve superior inference accuracy and speed compared to existing methods, but also enable dynamic prior updates without retraining while maintaining conjugacy—a unique capability in the field. Furthermore, DTs show competitive inference performance in terms of expected posterior NLL on challenging tasks, while maintaining the computational efficiency needed for practical applications.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. Yes
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. Not Applicable
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. Yes
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. Yes
    2. (b)
      Complete proofs of all theoretical results. Yes
    3. (c)
      Clear explanations of any assumptions. Yes
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). Yes
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). Yes
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). Yes
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). Yes
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. Not Applicable
    2. (b)
      The license information of the assets, if applicable. Not Applicable
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. /Not Applicable
    4. (d)
      Information about consent from data providers/curators. Not Applicable
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. Not Applicable
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. Not Applicable
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. Not Applicable
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. Not Applicable
Supplementary Materials
@@ Appendix B General Architecture Details
A standard transformer decoder setup is used, always consisting of 6 transformer decoder units, typically with a latent space size of 64, an MLP hidden layer size of 2048, and 8 attention heads. Layer norm was also chosen for its numerous benefits (Ziomek et al., 2025).
@@
For ACE, we also tried to use an similar setup as possible; for this region we used the GMM heads and used the same number of components and sample space transformations as in DTs. Additionally, since we are only interested in inference of certain variables in the problem, we did not randomise the variable with respect to which NLL loss is computed and instead kept it fixed to the variable of interest. We provide prior hyperparameters as additional latent variables on each problem.
For TabPFNv2, we simply used the pretrained foundational model provided by the authors and conditioned it on 10,000 samples (which is the maximum recommended by the authors). We provided prior hyperparameters as additional observations the model was conditioned on.
@@ Appendix C General Training Details
Table 4: Training hyperparameters and statistics for all experiments. An epoch always consists of 100 batches. Hyperparameters were obtained by limited manual tuning, adjusted until training was stable in all cases. For SVI, the total samples reported account for batching, so the number of samples used by each SVI problem will be a factor of the number of test problems lesser. Parameters are listed vertically, corresponding to DTs (top), PFNs (second from top), SVI (third from top) and ACE (bottom). For Experiment 4.1, only DT-5 is reported but training hyperparameters are identical for DT-2. For experiment 4.3 only DT is reported. For experiment 4.2.1, two PFNs are trained (one for each dimension, as PFNs do not trivially support multivariate distributions), so two figures are quoted when necessary (PPD + hyperposterior). Model sizes are not provided for VI as this is negligible.
@@
For the meta prior in this experiment, we assign a narrow, uniform prior over the PPD’s prior such that it is marginally everywhere a standard normal. We assign an inverse-gamma prior over lengthscale, whose hyperparameters are sampled from a uniform meta-prior, which is over $[100,105]$ for concentration and $[100,700]$ for rate.
@@
The sampling procedure for this experiment, using 5 observations (each observation being 5-dim input and 1-dim function value), was as follows:
@@
    Sample $x\sim\text{Uniform}(0,5)^{5}$
@@
    Sample $X\sim\text{Uniform}(0,5)^{5\times 5}$
@@
This experiment is a probabilistic adaptation of a well-known two-level quantum system problem (Nielsen & Chuang, 2010; Schorling et al., 2025). We adopt a beta prior, with an $\text{InverseGamma}(4,6)$ meta-prior over each parameter.